Team activity: Today we worked on rover calibration and software tuning in the morning in preparation for a system test in the afternoon. We prepped K10 Black for operations, mounted the GPR instrument, inflated the tires, and calibrated steering.
Matt Deans also spent a few minutes on a live EPO webcast going out to Challenger Centers with Matt Reyes and Leroy Chiao. He answered a few questions about the rovers and about the experiment we are doing.
Rover activity: K10 Black completed its shakedown test. We used the same plan that we did for Red yesterday just to keep thing simple and since we had some idea of where the rover would go.
Distance driven: K10 Black drove 610 meters today carrying out a survey. K10 Red stayed close to the rover shed for some low level testing and network testing. Remote operations from Johnson Space Center tested monitoring the rover activity from Houston.
When the rovers returned to the rover shed, Kimmick sauntered over and plopped down right in front of them and stood guard for a while.
Weather: Today it only got up to about 8 C, although it didn't feel too cold when standing in the sun. Some clouds but still another great day.
Quote of the Day: "You are all geeks," Pascal to the bunch of people still working in the Core when he was off to bed at 1:30 AM. Some of us were writing code, some were preparing bacteria samples, some of us were sorting photos and writing blog entries...